OpenStudy (anonymous):

Find the matrix that represents a skew dilation of R2 along the line y = 2x with scale factor 3, and the line y=0.5x with scale factor -3.

OpenStudy (anonymous):

So i have Ax=3x and Ax=-3x (A-3)x=0 (A+3)x=0

OpenStudy (anonymous):

when (A-3)x=0 x= (1,2) when (A+3)x=0 x= (2,1)